Raffles Bank pays 4% simple interest on its deposit account, whereas Bugis Bank pays interest on its deposit account compounded monthly. Formulate the quoted and effective interest rates that Bugis Bank should set if it wants to match Raffles Bank, assuming a 5-years horizon period


1
Expert's answer
2021-10-06T09:43:02-0400

"Effective\space interest\space rate=(1+interest \space rate×n)^{(\frac{1}{n})}-1\\=(1+4\%×5)^{(\frac{1}{5})}-1\\=3.7137\%\\\space \\

Quoted \space interest \space rate=((1+effective\space interest \space rate)^{(\frac{1}{12})}-1)×12\\=((1+3.7137\%)^{(\frac{1}{12})}-1)×12\\=3.6519\%"